10[(t+2)]^2]^2-17(t+2)^2+3=0 solve the equation.
Answer by stanbon(75887)

10[(t+2)]^2]^2-17(t+2)^2+3=0 solve the equation.
----
Let w = (t+2)^2
-------------
10w^2 - 17w + 3 = 0
Factor:
(2w-3)(5w-1) = 0
w = 3/2 or w = 1/5
-------------------------------------
(t+2)^2 = 3/2 or (t+2)^2 = 1/5
----
t+2 = sqrt(3/2) or t+2 = -sqrt(3/2) or t+2 sqrt(1/5) or t=2 = -sqrt(1/5)
----
etc..
Solve for "t":
Cheers,
Stan H.